# Self-Guided Open-Vocabulary Semantic Segmentation

### 07 Dec 2023 v1

@@ 4.4 Qualitative Analysis
Fig. 5 and App. B show qualitative results on the three datasets. Self-SEG is able to exhaustively identify relevant object classes present in the image while segmenting them accurately. Its ability to caption with specific class names such as hawk or treadmill showcases the model’s capabilities for OVS in the wild.
@@ 5 Conclusion
In this work we introduced Self-SEG, a novel method which leverages a vision-language model to automatically identify relevant object categories and segment them. In addition, we propose LOVE, a new evaluation method which employs an LLM to map from generated class names to ground-truth labels. Our method performs remarkably, achieving state-of-the-art performance on the zero label setting, while being competitive with open-vocabulary models which require ground-truth labels.
@@
Appendices
Appendix A LLM-based Open-Vocabulary Evaluator
@@
This appendix provides more insight into the pseudocode and text prompts of the LLM Mapper which maps open vocabulary classes to classes in the annotations of the dataset. The LLM Mapper is a key component of our LLM-based Open Vocabulary Evaluation Metric (LOVE) and enables the evaluation of predicted open vocabulary classes.
Pseudocode
The LLM Mapper is tasked with generating a mapping dictionary, which is used to change the predicted class index of each pixel to the corresponding class index in the dataset.
@@
     map_dict[skipped_key] $\leftarrow$ “background”
@@
LLM Prompts
The generateDialogs() function in the LLM Mapper takes a prompt template as one of its inputs. This prompt template is dependent on the dataset and is used to generate dialogs to query the LLM with. For Pascal VOC and CityScapes, we explicitly specify the list of categories, while for ADE20K we rely on the LLM’s knowledge of the dataset. Since ADE20K has many classes, dialogues get very long otherwise. This causes the LLM to run into memory issues when queried with dialogues in batches. We specify the prompt templates for each dataset as follows:
Pascal VOC and CityScapes
To which class in the list <dataset> is
<name> exclusively most similar to? If
<name> is not similar to any class in
the list or if the term describes stuff
instead of things, answer with
’background’. Reply in single quotation
marks with the class name that is part
of the list <dataset> and do not link
it to any other class name which is not
part  of the given list or ’background’.
ADE20K
To which class in the ADE20K dataset
exclusively most similar to? If <noun>
is not similar to any class in ADE20K
or if the term describes stuff instead
of things, answer with ’background’.
Reply in single quotation marks with
the class name that is part of ADE20K
and do not link it to any other class
name which is not part of the given
list or ’background’.
Appendix B Qualitative Results
This appendix provides additional qualitative results of Self-Seg. With individual sample outputs of the model, we show and discuss both its abilities and limitations.
Figure 6: Self-Seg is able to accurately classify and segment classes not present in the dataset (here Pascal VOC), such as the ox.
Figure 7: Interesting properties of objects are found with Blip-Cluster-Caption, such as reclining by the cat in this image.
Figure 8: Generated nouns are occasionally hyponyms or hypernyms of each other, causing the classes to “compete” for the final segmentation. Future work could possibly address this by checking for such relations between nouns and correcting for them.
Figure 9: Self-guidance occasionally yields a false positive noun, here man, which is then incorrectly used and segmented by X-Decoder.
Figure 10: The generated nouns are occasionally too generic (here animal and afar). The predicted mask is accurate nevertheless.
Figure 11: At times, an object or area is classified with one of its properties rather than the class name itself.
Figure 12: ADE20K has many classes, but badminton was not one of them. Despite being very small, it is still found and segmented.
Figure 13: Even with many objects present, Self-Seg maintains good performance in spotting and segmenting the objects.
Figure 14: Self-Seg can identify and segment an activity present in the image. Shown here is skiing with a white segmentation mask around the feet of the skier.
Figure 15: While in CityScapes all car types are considered equal, Self-Seg often provides more specific descriptions, such as the SUV in this image.
Figure 16: Self-Seg’s freedom in captioning the image with an open-vocabulary understanding of the world becomes evident in this example. Traffic lights in this scene are classified and segmented as signal.
Figure 17: Despite the visual similarities between a moped and motorcycle, Self-Seg correctly considers both. It is able to distinguish between the two vehicles and segment them accurately. A supervised model would not be able to generalize to moped, as it is not one of the annotated classes in CityScapes.

### 20 Mar 2024 v2

@@ 4.4 Qualitative Analysis
Fig. 5 displays qualitative results on the four datasets, where AutoSeg demonstrates its capability to comprehensively detect and accurately segment relevant object categories within the images. Remarkably, it predicts segmentation masks for specific class names such as hawk, piano, wildflower, officer or coke - all categories not listed in the annotations of their respective datasets. Moreover, in certain instances, AutoSeg successfully captures additional contextual details, such as graduation or exercise, illustrating the model’s capabilities at recognizing objects in scenes in a genuinely open-ended manner. Additional results, as well as failure cases, are available in the supplementary material.
@@ 5 Conclusion
In this work we introduced AutoSeg, a novel method which leverages a vision-language model to automatically identify relevant target classes and segment them - a task we termed Auto-Vocabulary Semantic Segmentation. Additionally, we proposed LAVE, a new evaluation method which employs an LLM to map open-ended class names to ground-truth labels. AutoSeg demonstrates remarkable open-ended recognition capabilities, achieving state-of-the-art performance on the zero label setting, while being competitive with open-vocabulary segmentation models which require ground-truth label specification.
@@
Supplementary Material
Appendix A. LLM-based Auto-Vocabulary Evaluator
@@
This appendix provides more insight into the pseudocode and text prompts of the LLM Mapper which maps auto-vocabulary classes to the fixed-vocabulary classes present in the annotations of each respective dataset. The LLM Mapper is a key component of our LLM-based Auto-Vocabulary Evaluation Metric (LAVE) and enables the evaluation of predicted auto-vocabulary classes on public benchmarks. By doings so, we enable comparison with with open-vocabulary methods, such as in Tab. 3, which predict a fixed-vocabulary for evaluation.
LLM Mapper Pseudocode
The LLM Mapper is tasked with creating a mapping dictionary which maps from auto-vocabulary classes to fixed-vocabulary classes. The resulting mapping dictionary can be used during inference to map pixel-level class predictions belonging to an automatically generated - out-of-vocabulary- class to fixed-vocabulary class indices. In Alg. 1, we detail the procedure to create the mapping dictionary.
LLM Mapping Prompts
The generateDialogs() function in the LLM Mapper takes a prompt template as one of its inputs. This prompt template is dependent on the dataset and is used to generate dialogs to query the LLM with. For PASCAL VOC and Cityscapes, we explicitly specify the list of categories given its short length with 20 classes each. For PASCAL Context and ADE20K, with its 459 and 849 classes respectively, we rely on the LLM’s knowledge of the datasets. Explicitly naming each object class from these datasets causes the LLM to run into memory issues when queried with long dialogues. We specify the prompt templates for each respective dataset as follows:
PASCAL VOC and Cityscapes
To which class in the list <dataset> is <name> exclusively most
similar to? If <name> is not similar to any class in the list or
if the term describes stuff instead of things, answer with
‘background’. Reply in single quotation marks with the class name
that is part of the list <dataset> and do not link it to any other
class name which is not part of the given list or ‘background’.
PASCAL Context and ADE20K
To which class in the <dataset> dataset is <name> exclusively most
similar to? If <name> is not similar to any class in <dataset> or
if the term describes stuff instead of things, answer with
‘background’. Reply in single quotation marks with the class name
that is part of <dataset> and do not link it to any other class
name which is not part of the given list or ‘background’.
where <noun> is the noun to map to the vocabulary and <dataset> is either ‘PASCAL-Context 459’ or ‘ADE20K’.
Appendix B. Alternative Class Generation Baselines
In Tab. 4, we provide additional comparisons with alternative methods to generate class names from images. In particular, we compare against an altered version of BLIP to caption the image at an object instance level with segments retrieved from the Segment Anything Model (SAM[12]. We use segment everything prompting, i.e. the model’s object proposal generation functionality. Furthermore, we compare against Large Language and Vision Assistant (LLaVa[18], generating classes with the prompt ‘Please name all the elements present in the image as a list of nouns’. We compare the mIoU performance on VOC, Context and Cityscapes and find that AutoSeg performs favorably over other approaches that are adjusted to segment with automatically generated classes. The performance gap is especially high for datasets with a small vocabulary, such as VOC and Cityscapes, possibly indicating that AutoSeg is better able in generating classes that are relevant for prediction while the other methods generate redundant classes which become false positive predictions. In contrast to both SAM and LLaVa, AutoSeg has an additional benefit of being independent of prompting. For this reason, our method does not require prompt engineering in order to obtain class names effective for accurate segmentation.
Table 4: Experiments with Alternative Class Generation Baselines (mIoU). Experiments with alternative methods to generate classes showcase that AutoSeg is favorable for AVS across all three datasets, with larger performance gaps on VOC and Cityscapes with a smaller vocabulary size.
Approach VOC [9] PC-459 [23] Cityscapes [6]
SAM + BLIP + X-Decoder 41.1 11.3 27.4
LLAVA + X-Decoder 56.7 11.4 23.4
AutoSeg 87.1 11.7 29.5
Appendix C. Visualization of Embeddings in BCC
In Fig. 6, we visualize intermediate 32x32 outputs of BCC to better understand what each individual step of Clustering, Alignment or Denoising achieves. Firstly, Fig. 6 reveals that clustering BLIP embeddings results grouping pixels into semantic areas. With $k=2$, the image is divided an area which includes the cows (in blue) and another for everything else (green), while for $k=8$, various other objects - such as the pole in purple - get included. At this stage, the embeddings appear noisy and inconsistent across all $k$, which is why the combination of all clusters is so important. The alignment output shows that certain objects from a higher $k$ embeddings are now included in those of lower $k$ embeddings, i.e. the aligned embedding captures a more extensive list of objects present in the image. Finally, to further reduce the noise in the embedding, we apply CRF and a majority filter. We see that the final embedding consists of three unique semantic areas, which are eventually fed to the BLIP text decoder for captioning.
Figure 6: Visualization of Intermediate BCC Outputs. Intermediate BCC outputs show how BLIP embeddings are enhanced with each individual step of Clustering, Alignment or Denoising.
Appendix D. Additional Qualitative Results
In Fig. 7,  8 and  9, we provide additional qualitative results of AutoSeg and discuss observed capabilities or limitations in the caption of each figure.
(a) Success Case. AutoSeg can infer detailed class names, such as courtroom, which is non-existent in the extensive vocabulary of the dataset.
(b) Success Case. The class badminton is not part of ADE20K. Despite taking up a very small area in the image, AutoSeg is still capable of segmenting it.
(c) Success Case. AutoSeg displays advanced scene understanding capabilities. Here, it recognizes the out-of-vocabulary class reflection, which requires the model to reason about other auto-generated classes and structural properties in the image.
(d) Success Case. The rare and out-of-vocabulary class name dinosaur is generated automatically and segmented by AutoSeg.. This is another example of the method’s efficacy on long-tailed classes.
(e) Failure Case. AutoSeg’s occasionally struggles with indoor scenes from specific buildings, such as a mosque in this image. Our model mistakenly regards it a temple.
(f) Failure Case. Here, AutoSeg predicts the zebra rug as a zebra as a result of noun parsing. After BCC has generated captions from the image, both zebra and rug get parsed into individual nouns. When fed to the segmentor, zebra eventually ends up being predicted as the more relevant class name.
Figure 7: Additional Qualitative Results on ADE20K.
(a) Success Case. Despite the visual similarities between a moped and motorcycle, AutoSeg correctly considers both and segments them accurately. A supervised model would fail to generalize to moped, as it is out of Cityscapes’ vocabulary.
(b) Success Case. While annotations in Cityscapes consider all car types to be equal, AutoSeg often generates more specific descriptions, such as the SUV in this image.
(c) Success Case. AutoSeg is able to provide class names which can be crucial for real-life scenarios, such as hazard in this example. The standard 20 classes in Cityscapes do not account for such objects descriptions.
(d) Failure Case. Occasionally, generated class names, such as street, dominate the search space of the segmentor and cause it to attend more on that class. This may result in oversegmentation and other class names, such as car, to be ignored.
Figure 8: Additional Qualitative Results on Cityscapes.
(a) Success Case. AutoSeg is able to accurately classify and segment classes not present in the dataset, such as the ox.
(b) Failure Case. Generated nouns are occasionally hyponyms or hypernyms of each other, causing the classes to “compete" for the final segmentation. Future work could possibly address this by checking for such relations between nouns and correcting for them.
Figure 9: Additional Qualitative Results on PASCAL.
@@ Appendix D. Additional Qualitative Results
     map_dict[skipped_key] $\leftarrow$ “background"